OpenStudy (anonymous):

the vector a is parallel to (1,1,1) and has length 3. find a.

OpenStudy (anonymous):

there are infinite solution for what you are asking. Can you please state your question in other terms?

OpenStudy (amistre64):

we need to find a scalar, if anything, for the given vector that will stretch it out to 3. |<1,1,1>| = sqrt( 1^2+1^2+1^2) = sqrt(3) N sqrt(3) = 3 N = 3/sqrt(3) = sqrt(3) <sqrt(3),sqrt(3),sqrt(3)> is one such vector that meets the criteria